please help me solve this!!! SUPER CONFUSED & this is on a test tomorrow that i CAN NOT FAIL!!!!
3x + 5 = 1 + 6 - 5x
????? i have no idea!!

Found 2 solutions by Alan3354, edjones:
Answer by Alan3354(69443)

3x + 5 = 1 + 6 -5x
-----------
On the RHS (right side) add the 1 and the 6.
Then add 5x to both sides
Then subtract 5 from both sides.
Then divide by 8
----------------
I won't be able to accompany you on your test, I have things to do.

Answer by edjones(8007)

3x + 5 = 1 + 6 - 5x
3x+5=7-5x
8x=2 add 5x-5 to each side
x=1/4 divide each side by 8.
.
Ed